OpenStudy (anonymous):

Eight more than the product of 5 and a number is equal to the number. Find the number.

ganeshie8 (ganeshie8):

call the number x

ganeshie8 (ganeshie8):

product of 5 and a number : 5x

ganeshie8 (ganeshie8):

Eight more : 5x + 8

ganeshie8 (ganeshie8):

it equals the original number x itself, so : 5x + 8 = x

ganeshie8 (ganeshie8):

you can solve for x ?

OpenStudy (anonymous):

thx i got this under control

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

5x - x = -8 4x = -8 x = -8/4 x = -2

ganeshie8 (ganeshie8):

\(\checkmark\)